# Cloud Storage Integrity at Scale:

### v2

@@ Abstract.
Merkle hash trees are the state-of-the-art method to protect the integrity of storage systems. However, using a hash tree can severely degrade performance, and prior works optimizing them have yet to yield a concrete understanding of the scalability of certain designs in the context of large-scale cloud storage systems. In this paper, we take a first-principles approach to analyzing hash tree performance for storage by introducing a definition of an optimal hash tree and a principled methodology for evaluating hash tree designs. We show that state-of-the-art designs are not scalable; they incur up to 40.1×40.1$\times$40.1 × slowdowns over an insecure baseline and deliver ¡50% of optimal performance across various experiments. We then exploit the characteristics of optimal hash trees to design Dynamic Hash Trees (DHTs), hash trees that can adapt to workload patterns on-the-fly, delivering ¿95% of optimal read and write performance and up to 4.2×4.2$\times$4.2 × speedups over the state-of-the art. Our novel methodology and DHT design provides a new foundation in the search for integrity mechanisms that can operate efficiently at scale.
@@ I. Introduction
An increasing number of attacks against cloud services has raised significant concerns about how to practically secure data when outsourcing storage resources. This has fueled significant investment and research into trusted cloud storage systems: storage systems that provide high assurance of the confidentiality and integrity of data in-memory and on-disk through hardware-based enforcement mechanisms (McKeen et al., 2013; che Tsai et al., 2017; Arnautov et al., 2016). To this end, Merkle hash trees have become the state-of-the-art method to protect the integrity of trusted cloud storage systems (Merkle, 1989; McKeen et al., 2013; che Tsai et al., 2017; Priebe et al., 2020; Android, 2023; Arasu et al., 2021).
[…]
Several works have examined various hash tree optimizations, mostly in the context of secure memory systems (Gassend et al., 2003; Feng et al., 2021), but also recently for storage (Merkle, 1989; McKeen et al., 2013; che Tsai et al., 2017; Priebe et al., 2020; Android, 2023; Arasu et al., 2021). We focus specifically on storage, which is subject to vastly different workload characteristics, capacities, and assumptions about cache memory size. While state-of-the-art works optimizing hash trees have claimed to be scalable, we observe that they are not—incurring up to 40.1$\times$ slowdowns under realistic workloads. We attribute this inconsistency to the fact that prior works have yet to yield a principled framework within which to evaluate hash tree performance.
@@
In this paper, we take a first-principles approach to analyzing hash tree performance in the context of cloud block storage. First, we observe that prior works lack a well-defined performance target: it remains unclear how to evaluate whether a hash tree design is performing optimally (Keeton et al., 2007), or even what performance is practically achievable, in a given system. Understanding optimal performance is important to understanding scalability limits. We ask the fundamental question: How can we model an optimal hash tree for cloud block storage? We show that the problem of finding an optimal hash tree can be reduced to the problem of finding an optimal prefix tree in the context of lossless data compression (Huffman, 1952). More specifically, by constructing a hash tree as an optimal prefix code, we can produce a hash tree that achieves optimal throughput under a given workload profile.
Next, we observe the lack of a principled methodology to qualitatively evaluate how well a design performs across the various parameters that characterize cloud block storage deployments, including capacity, read/write ratio, cache size, and workload type (Arasu et al., 2021; Sinha and Christodorescu, 2018). For example, a hash tree may appear to have high overheads, but actually be performing optimally, which suggests that complimentary optimizations (e.g., parallelization) may be the only way to break the performance ceiling. Towards this, we introduce a new method to instantiate an optimal hash tree in an offline setting (i.e., with a priori knowledge of a workload trace) and evaluate how well a design performs with respect to the optimal.
@@
Building on our observations of optimal trees, we then explore the design of an online solution that does not assume a priori knowledge of workload characteristics, but can learn and adapt to workload patterns on-the-fly. It is know that real-world workloads are characterized by skewed access patterns (i.e., with high reference locality, where a small number of blocks are accessed much more frequently than others) across all layers of the memory hierarchy (Li et al., 2023; Wang et al., 2023; Cooper et al., 2010; Yang and Zhu, 2016; Arasu et al., 2021). In an offline setting, this manifests as optimal hash trees often being far from balanced. Thus, frequently accessed blocks have shorter verification/update paths in the tree than infrequently accessed blocks. Towards this, we introduce a novel dynamic, unbalanced hash tree design called Dynamic Hash Trees (DHTs). DHTs are based on the splay trees commonly used in garbage collection and IP routing lookup (Sleator and Tarjan, 1985), and they self-adjust to remain adaptive to workload characteristics observed at runtime.
@@
We implemented DHTs and the state-of-the-art designs and evaluated them in a real cloud setting with AWS EC2 instances and EBS block devices. Our experiments under realistic workload patterns and across various system configurations show that: (1) state-of-the-art static, balanced trees (Merkle, 1989; Feng et al., 2021; Gassend et al., 2003; che Tsai et al., 2017; Freij et al., 2021) are not scalable, incurring up to 40.1$\times$ slowdowns over an insecure baseline and delivering less than ¡50% of optimal throughput on average; and (2) DHTs can capitalize on skewed access patterns, delivering ¿95% of optimal throughput and up to 4.2$\times$ speedups on read and write performance over the state-of-the-art.
@@
We conclude that for cloud block storage, balanced trees are ill-suited as the base construction of a hash tree, and DHTs are a preferable alternative. DHTs are an existence proof that online approaches are feasible to implement and even simple heuristics can significantly improve performance over the state-of-the-art. Our methodology and DHT design provides a new foundation in the search for integrity mechanisms that can operate efficiently at scale. Our code is plug-and-play into standard Linux systems and is open-sourced at [Anonymized link].
@@ II. Background
Cloud Block Storage. Block storage systems are a backbone of modern cloud infrastructure. As shown in Figure 1, they sit at the lowest level in the software stack and serve as the primary storage interface for various higher-level software. They are composed of three key components: clients, block layer, and cloud storage devices. Clients are any software requiring persistent storage, such as remote user clients of personal cloud drives, virtual machines, cloud-hosted applications, etc. The client’s block layer implements protocols to communicate with local or remote cloud storage devices (also called disks or volumes) and exposes a simple block read/write interface to clients. Widely used cloud block storage services include AWS Elastic Block Store (AWS, 2023), Google Cloud Persistent Disks (Cloud, 2023), and Microsoft Azure Managed Disks (Azure, 2023)—none of which provides cryptographic integrity protection to customers by default.
Figure 1. Our system model captures common deployment scenarios for cloud block storage. The block layer (and client code) runs inside a trusted environment, while cloud storage devices managed by the public cloud provider are untrusted.
@@
Merkle Hash Trees. Merkle hash trees have become the state-of-the-art method to protect the integrity of arbitrary datasets (Merkle, 1989; Gassend et al., 2003; McKeen et al., 2013; che Tsai et al., 2017; Priebe et al., 2020). As shown in Figure 2, a Merkle hash tree (or more simply, a hash tree) is typically a balanced binary tree, with each node in the tree containing a hash value. We note that higher-degree trees (e.g., 64-ary) have been considered in other contexts such as secure memory systems (Taassori et al., 2018); we defer discussion of these designs to Section VII. A leaf node contains the hash (MAC) of a data block (and a cipher IV when encrypting data), and an internal node contains the hash of the concatenation of the hashes of its two children. Internal node hashes are iteratively computed from leaf to root, and the root hash authenticates the current contents of the storage device and is typically stored in a secure location (e.g., in a TPM on the client machine (Perez et al., 2006)). All other nodes in the tree are stored on disk alongside the data. The number of leaf nodes in the tree $n$ is equal to the number of blocks on the storage device, and the total number of tree nodes is $12n-$ 1.
@@
There are two primitive operations on a hash tree: verification and update. When a block is read, it must be verified against the root hash. The client’s block layer first fetches the (encrypted) block data, MAC, and cipher IV from disk. It checks that the retrieved MAC is consistent with the retrieved block data by rehashing the data and comparing. It then fetches the proof of authenticity, a set of auxiliary hashes along the path from the accessed leaf to the root (see nodes highlighted in blue). The retrieved MAC is inserted into the tree at the appropriate leaf position, and parent hashes are iteratively computed along the authentication path using the auxiliary hashes (see red arrows). The computed root hash is then compared against the known root hash. If the two hashes match, verification succeeds. Block updates are handled similarly, with the update committed by saving the new root hash to the secure location.
@@
Caching hashes in a fast memory is also a standard hash tree optimization (Gassend et al., 2003; Arasu et al., 2021). Caching provides two advantages: (1) it reduces I/O costs associated with fetching hashes during a verification or update, and (2) it enables early returns when fetching and verifying nodes (during either a verification or update), because cached nodes are already authenticated.
Figure 2. A Merkle hash tree protects against data corruption, relocation, and replay attacks with a tree of hashes. The client’s block layer uses a Merkle hash tree to authenticate data when it is read from or written to storage.
@@ III. Security Model
Trust Model. We assume the clients and block layer are trusted, while cloud storage devices managed by the public cloud provider are untrusted. This trust model broadly captures practical deployment scenarios of trusted storage (Bocchi et al., 2015; Kallahalla et al., 2003; Srinivasan et al., 2015; Arnautov et al., 2016). For example, for trusted cloud drives, the block layer and all other client code would run in a trusted environment, the remote user’s machine, and send read/write requests to storage over a network. For trusted VMs or cloud-hosted applications, the block layer and all other client code would run inside of a trusted execution environment (McKeen et al., 2013) on a cloud server, similarly sending read/write requests to local or remote storage (che Tsai et al., 2017; Priebe et al., 2020). In either case, the block layer runs underneath the application code, within the trust boundary, and bridges the application to untrusted block storage.
Threat Model. Attacks on cloud block storage may come from either a malicious cloud provider, co-tenant, or another authorized client of the block storage service (che Tsai et al., 2017; Feng et al., 2021). We focus specifically on attacks causing integrity violations: any block written to or read from disk may be tampered with. For example, an attacker may corrupt data blocks before they are delivered up to the block layer. They may perform relocation attacks, where blocks are deliberately read from or written to incorrect block addresses. Finally, they may replay an old version of a block to the block layer.
Security Requirements. Three integrity properties are required of trusted cloud block storage: correctness (protects against corruptions), spatial uniqueness (protects against relocations), and temporal uniqueness (protects against replays) (Avanzi et al., 2022). Merkle hash trees are the standard method to ensure all three properties, and in particular, replay protection (which per-block hashes cannot ensure). Their use is becoming increasingly important in practice for both storage and memory (Gassend et al., 2003; Arasu et al., 2021; Angel et al., 2023; Matetic et al., 2017). For example, hash trees can prevent an adversary from reverting (i.e., replaying) system/application software to a known vulnerable state: the root hash reflects the current version of the storage device, so a replay attack would require changing the root hash, which is stored in a secure location and is out of control of the attacker.
We note that to achieve better performance, some prior works have loosened security requirements by permitting lazy verification or splitting the hash tree into multiple independent trees (Arasu et al., 2021; Freij et al., 2021). However, lazy verification introduces a window of vulnerability, during which an attacker can violate one or more integrity properties. Splitting the hash tree also precludes having global correctness for any data within a single security domain. We therefore do not consider lazy verification or splitting in our analysis.
IV. Optimal Hash Trees
Prior works have explored various hash tree optimizations (Gassend et al., 2003; Feng et al., 2021; Sinha and Christodorescu, 2018; Arasu et al., 2021), but have yet to concretely characterize what techniques perform well (or best) and when (Keeton et al., 2007). We address this problem by introducing a novel definition of an optimal hash tree. Having an optimal tree serves two purposes: it establishes an upper bound on performance (Section V), and it helps to understand what characteristics of the tree structure are correlated with high-performance, which will help inform new designs (Section VI).
IV-A. Optimal Definition
Overheads from storage-level mechanisms are largely attributable to two things: compute costs and I/O costs. For hash trees, this means that if cache performance is poor, then I/O costs may be the bottleneck. In contrast, if cache hit rates are high, then performance is directly related to how quickly hashes can be computed to complete a verification or update—that is, how efficient the base structure of the hash tree is. Prior works have identified that even small caches are very efficient, but as we will show in Section VII, state-of-the-art static, balanced hash trees still cannot scale to large capacities; this is largely due to an increased tree size and thus increased hashing costs. This motivates our search for a more efficient tree structure.

### v3

# On Scalable Integrity Checking for

@@ Abstract.
Merkle hash trees are the standard method to protect the integrity and freshness of stored data. However, hash trees introduce additional compute and I/O costs on the I/O critical path, and prior efforts have not fully characterized these costs. In this paper, we quantify performance overheads of storage-level hash trees in realistic settings. We then design an optimized tree structure called Dynamic Merkle Trees (DMTs) based on an analysis of root causes of overheads. DMTs exploit patterns in workloads to deliver up to a $\times2$ throughput and latency improvement over the state of the art. Our novel approach provides a promising new direction to achieve integrity guarantees in storage efficiently and at scale.
@@ I. Introduction
An increasing number of attacks against cloud services has fueled significant investment and research into trusted cloud storage systems: systems that provide high assurance of the confidentiality and integrity of data stored in-memory and on-disk through hardware-based access-controls and cryptographic proof systems (tsai_graphene-sgx_nodate, ; priebe_enclavedb_2018, ; arasu2021fastver, ). To this end, using a Merkle hash tree (merkle1989certified, ) has become the state-of-the-art method to protect the integrity and freshness of both volatile (taassori2018vault, ; feng2021scalable, ) and persistent (arasu2021fastver, ; chakraborti2017dm, ) storage. An exemplar use case is protecting disks attached to confidential virtual machines (aws_sev_snp, ; gcp_confidential_vm, ; azure_confidential_vm, ).
However, hash trees introduce additional compute (hashing) and I/O (metadata fetching) costs on the I/O critical path, which can severely degrade performance. For example, consider that a 1 TB disk contains $\approx$268 M 4 KB blocks. A typical balanced, binary hash tree over the disk blocks would have a height of 28, requiring (at least) 28 hashes to be computed on every read or write. The total cost of fetching metadata and verifying/updating hashes can exceed several hundred $\mu s$, dwarfing the baseline latency of performing a data access on a high-performance storage device (which can be $<60~{}\mu s$).
Prior works have studied this phenomenon, primarily in the context of secure volatile memory (taassori2018vault, ; yan2006improving, ; gassend2003caches, ; feng2021scalable, ). However, their performance implications in the context of (cloud) block storage at large, and low-latency storage devices in particular, remain largely unknown. The key difference is that storage devices are subject to vastly different workload characteristics, capacities, and cache behaviors than memory devices.
@@
In this paper, we take a first-principles approach to analyzing hash tree performance in the context of cloud block storage. First, we demonstrate that state-of-the-art hash tree designs incur significant overheads and fail to reliably scale to large disk capacities. We then demonstrate that hashing (CPU) costs are the primary performance bottleneck. Next, we address the challenge of reducing hashing costs by posing the fundamental question: How can we model an optimal hash tree for cloud block storage? We show that the problem of finding an optimal hash tree can be reduced to the problem of finding an optimal prefix tree in the context of lossless data compression (huffman1952method, ). More specifically, by constructing a hash tree as an optimal prefix code, we can produce a hash tree that achieves optimal throughput under a known workload profile.
@@
Building on our observations of optimal trees, we then develop an online solution that can approximate an optimal tree (without a priori knowledge) by learning and adapting to workload patterns on-the-fly. It is known that real-world workloads are characterized by skewed access patterns (i.e., where a small number of blocks are accessed much more frequently than others) across all layers of the memory hierarchy (li2023depth, ; wang2023locality, ; cooper2010benchmarking, ; yang2016write, ; arasu2021fastver, ). In an offline setting, this manifests as optimal hash trees often being far from balanced—where frequently accessed blocks have shorter verification/update paths in the tree than infrequently accessed blocks. Towards this, we introduce a novel dynamic, unbalanced hash tree design called Dynamic Merkle Trees (DMTs). DMTs are based on the splay trees commonly used in garbage collection and IP routing (sleator1985self, ), and they self-adjust at runtime to reduce hashing costs for frequently accessed data.
@@
We implemented DMTs and evaluated them in a real cloud setting with AWS EC2 instances and NVMe devices. We performed a broad performance analysis across a range of system and workload settings, parameterized by disk capacity, read/write ratio, I/O size, etc. Using a set of Zipfian workloads, an Alibaba dataset, and a Filebench OLTP workload, we show that the static nature of state-of-the-art approaches becomes prohibitive: they deliver less than 50% of optimal throughput on average across all experiments. In contrast, DMTs capitalize on skewed access patterns, delivering ¿85% of optimal throughput and up to a $\times2$ throughput and latency improvement over the state of the art.
@@
We conclude that for cloud block storage, balanced trees are ill-suited as the base construction of a hash tree, and DMTs are a preferable alternative. DMTs provide a new foundation in the search for integrity mechanisms that operate efficiently at scale. Our code is plug-and-play into Linux and open-sourced at https://github.com/MadSP-McDaniel/dmt.
@@ II. Background
Cloud Block Storage. Block storage is a backbone of modern public cloud infrastructure (aws-ebs, ; gc-pdisk, ; azure-mdisk, ). While there are various deployment models for cloud applications and storage, we consider a standard Infrastructure-as-a-Service (IaaS) deployment where an application runs inside of a guest VM and reads and writes to a fast, local NVMe disk attached to the VM (Figure 1). The application may be end-user facing (e.g., a web server) or the last hop in a networked storage system (e.g., a file server for other cloud-hosted applications).
Figure 1. We assume that VM memory contents are trusted and cloud storage devices are untrusted; VM memory can be protected with trusted execution primitives (aws_sev_snp, ).
Merkle Hash Trees. Merkle hash trees are the state-of-the-art method to protect the integrity and freshness of arbitrary datasets—largely due to their proven theoretical efficiency (merkle1989certified, ; gassend2003caches, ; mckeen_innovative_2013, ; tsai_graphene-sgx_nodate, ; priebe_sgx-lkl_2020, ). They have played a pivotal role in ensuring boot disk integrity with Linux dm-verity (android-dm-verity, )—where they are implemented as a custom device driver that intercepts I/Os and implements the hash tree logic.
@@
As shown in Figure 2, a Merkle hash tree (or more simply, a hash tree) is typically a balanced binary tree, with each node in the tree containing a hash value. A leaf node contains the hash (MAC) of a data block (and a cipher IV when encrypting data), and an internal node contains the hash of the concatenation of the hashes of its two children. Internal node hashes are iteratively computed from leaf to root. The root hash authenticates the current contents of the storage device and is typically stored in a secure location (e.g., a persistent on-chip register or a TPM (perez2006vtpm, ; taassori2018vault, )). All other nodes in the tree are stored on disk alongside the data. The number of leaf nodes in the tree $n$ is equal to the number of blocks on the storage device, and the total number of tree nodes is $12n-$ 1.
Figure 2. A Merkle hash tree protects the integrity and freshness of data read from/written to a storage device.
@@
There are two primitive operations on a hash tree: verification and update. When a block is read, it must be verified against the root hash. The client’s block layer first fetches the (encrypted) block data, MAC, and cipher IV from disk. It checks that the retrieved MAC is consistent with the retrieved block data by rehashing the data and comparing. It then fetches the proof of authenticity, a set of sibling hashes along the path from the accessed leaf to the root (see nodes highlighted in blue). The retrieved MAC is inserted into the tree at the appropriate leaf position, and parent hashes are iteratively computed along the authentication path using the sibling hashes (see red arrows). The computed root hash is then compared against the known root hash. If the two hashes match, verification succeeds. When a block is written, a new hash must be computed (H0) and the hash tree updated (H0 ancestors). Updates are handled similar to verification, with a new root hash computed and saved to the secure location.
@@
Caching hashes in secure memory (i.e., in a protected memory region) is also a standard hash tree optimization (gassend2003caches, ; arasu2021fastver, ). Caching reduces I/O costs associated with fetching hashes during a verification or update. It also enables early returns when fetching and verifying hashes; cached hashes were already authenticated, so any hash that has been tampered with will result in a failed check at some level in the tree.
@@ III. Security Model
Data-only attacks—attacks based on maliciously crafted data rather than control flow hijacking—have been recently shown to present a significant threat to modern applications (johannesmeyer2024practical, ; kurmus2017random, ; galanou2023trustworthy, ; bohling2020subverting, ; hu2016data, ). Any attack that can be launched from (untrusted) storage would fundamentally upend the security guarantees provided by the rest of the system. Below we describe these attacks and outline security requirements to mitigate them.
Trust Model. We assume all VM contents (code and data) are trusted and the storage devices are untrusted. VM memory contents can be protected with hardware-based isolation primitives such as AMD SEV-SNP (aws_sev_snp, ). The trusted and untrusted components therefore have a simple block read/write interface (Figure 1). This models untrusted disks attached to, for example, confidential virtual machines (aws_sev_snp, ; gcp_confidential_vm, ; azure_confidential_vm, ).
Threat Model. We consider a privileged attacker who has access to the hypervisor or storage backbone in a public cloud datacenter (tsai_graphene-sgx_nodate, ; arnautov_scone_nodate, ). This could be a malicious co-tenant who was able to escalate privilege, or a malicious cloud administrator. The attacker has the ability to access, corrupt, swap, drop, record, inject, or replay any data across the storage backbone.
Example attacks. Armed with the capability to inject arbitrary data into the storage interface, the attacker could replay old data to the VM (kurmus2017random, ; johannesmeyer2024practical, ). Data would bubble up the call stack and either cause the VM to deliver old data to applications, or cause an outdated version of a binary to be read from disk and executed (johannesmeyer2024practical, ; galanou2023trustworthy, ; bohling2020subverting, ). Similarly, consider an ext4 file system formatted on top of the disk. An attacker could arbitrarily replay inode table blocks and cause the VM OS to recognize an invalid set of permissions on a file, enabling unauthorized access to the file. Checksums or keyed hashes alone cannot prevent these data-only attacks: the received data would still pass verification.
Security Requirements. Ensuring the safety of user data and correct execution of applications therefore requires three security properties for storage: authenticity, uniqueness, and freshness (avanzi2022cryptographic, ). Keyed encryption and MACs can ensure confidentiality, authenticity (prevents corruptions), and uniqueness (prevents relocation attacks). Merkle hash trees then ensure data freshness (gassend2003caches, ; arasu2021fastver, ; angel2023nimble, ; matetic2017rote, ): the root hash reflects the current version of the storage device, so a replay attack would require changing the root hash, which is stored in a secure location and is out of control of the attacker  [^1] To improve performance, some prior works have loosened security requirements by permitting lazy verification (arasu2021fastver, ). However, this violates freshness guarantees; we therefore do not consider lazy verification in our analysis..
IV. Motivation
Though hash trees have played a pivotal role in ensuring boot disk integrity with Linux dm-verity (android-dm-verity, ), their performance implications in the context of (cloud) block storage at large, and low-latency storage devices in particular, are largely unknown. In fact, prior works have identified that hash tree overheads can severely degrade performance for secure memory systems, and optimizations abound (taassori2018vault, ). This raises the natural question of whether storage-level hash trees observe similar costs. Our goal in this paper is to quantify this effect and design optimizations to reduce overheads if so.
Figure 3. This graph shows how throughput decreases w.r.t. capacity under an exemplar setup and workload. Experiment parameters: Workload: Zipf(2.5), Read ratio: 1%, I/O size: 32 KB, Cache size: 10%.
Figure 4. CPU vs. I/O time during the driver write routine. Same experiment parameters as above.
Scalability Problem. We begin with a motivating experiment in Figure 3, which demonstrates the performance of the state-of-the-art hash tree design used by dm-verity—a balanced, binary tree. The graph shows how throughput changes as disk capacity increases. We defer implementation and experiment setup details to Section VII-A, but note that the hash tree is implemented in a block device driver that wraps a lower-level driver, and is exposed as a regular device to file systems or other applications as /dev/XXX.
The graph shows that throughput decreases w.r.t. capacity. This is due to the tree size (height) increasing logarithmically with capacity, which is reflected in logarithmically increasing slowdowns. At 16 MB capacity, the hash tree incurs nearly a 60% throughput loss over the Encryption/no integrity baseline. At 4 TB capacity, throughput loss increases to 75%. Note that the workload shape is immaterial here; the same overheads are always observed because of the tree structure (i.e., all accesses will require computing the same number of hashes).
Read-heavy workloads do not pose significant challenges, because the (small) hash cache is very efficient (hit rate ¿99%), and verifies benefit from early exits when they hit a cached hash. The problem is how to efficiently handle writes: under write-heavy workloads, hash tree overheads are prohibitive at small and large capacities, undermining the performance capability of the fast NVMe device [^2] Note that in some cases the page cache can hide disk-level performance impacts, but the performance of large-scale applications will depend heavily on the underlying disk performance (due to writeback contention and throttling under heavy memory pressure (corbet2010block, )). Our focus is thus on analyzing and optimizing the storage layer to ensure that the performance capabilities of fast NVMe devices are not undermined by hash tree operations..
Root Cause Analysis. Figure 4 shows the latency breakdown during the device driver write routine. As expected, for a 32 KB I/O the time spent pushing data out to disk (data I/O) is approximately $60~{}\mu s$. The remaining time in the write routine is spent fetching/writing hashes to disk (metadata I/O) and performing hash updates (computing the new block hash and executing the hash tree update). Metadata I/O is negligible because the hash cache is very efficient. The majority of time is therefore attributed to managing the hash tree [^3] Note that our focus is on NVMe SSDs; HDDs have a different performance profile that we are not optimizing for. With HDDs, the data access time dominates hashing time, so hash tree overheads can be relatively negligible..
To understand why this occurs, Figure 5 shows the latency to compute a SHA256 hash (the standard hash function used in Merkle hash trees) vs. data size on a 2.9GHz Intel Xeon Platinum 8375C, a 3rd Generation Intel Xeon Scalable processor supporting AES and SHA instruction-set extensions to accelerate cryptographic operations. We observe that it takes approximately $490~{}ns$ to compute the hash of 64 B of data. We also measure the latency to encrypt and generate the MAC for a 4 KB block with AES GCM to be approximately $2~{}\mu s$.
Figure 5. This graph shows the latency of computing SHA256 hashes on a modern processor with hardware acceleration for cryptographic functions. The annotations highlight the input data size to the hash function at different tree arities.
Figure 4 shows that at 1 GB capacity, approximately $150~{}\mu s$ is spent managing the hash tree. Consider that a 1 GB disk has 262,144 4 KB blocks and thus a height of 18, requiring one SHA256 computation per level. Further, with 4 KB disk blocks, executing a 32 KB write I/O would require $32768/4096=8$ hash tree updates executed sequentially—best-known methods still rely on a global tree lock to serialize tree updates. This amounts to $150~{}\mu s/8=18.75~{}\mu s$ spent encrypting data, generating the MAC, and updating the hash tree. Thus, we have $18.75-2=16.75~{}\mu s$ time spent doing the actual hash tree update, and $16.75/18=0.93~{}\mu s$ total time spent doing work at each level in the tree. Most of this time is spent computing the node hash, with the remaining work being cache lookups and buffer copying.
Note that with even faster devices in the future (with single-digit microsecond access latencies), the proportion of time spent hashing vs. doing data I/O will grow substantially.
Optimized Tree Structures. Fundamentally, this means that time spent hashing (CPU costs) is the bottleneck. Section VII will show that caching and parallelization only help to an extent. What is needed is a structurally more efficient tree.
Prior works optimizing hash trees for memory have largely converged on the idea that high-degree (e.g., 64-ary) trees are the solution to eliminate overheads (taassori2018vault, ). The intuition is that by increasing tree fanout, one can decrease tree height and thus the number of hashes that must be computed per read or write. However, Figure 6 shows that high-degree trees are actually a suboptimal design choice. We compute the expected hashing costs based on the hashing latency and the height of the tree observed under a given arity for 1 GB capacity (e.g., 64-ary trees have height 3). While tree height is decreased, the graph shows that increased fanout results in high-degree trees incurring the highest expected hashing costs due to hashing more content, and ultimately lower performance. We will demonstrate in Section VII that high-degree trees fail to reliably scale, and binary trees perform best. We therefore seek a better way to maneuver binary trees to reduce overheads.
Figure 6. We calculate the expected hashing costs for a 32 KB write I/O based on the tree height under different tree arities, given the measured SHA256 latencies for each arity in Figure 5. The graph shows that low-degree trees should have lower hashing costs than high-degree trees.
V. Optimal Hash Trees
We approach this problem by asking the fundamental question: Is there an optimal tree structure? (keeton2007don, ). Having a definition of an optimal tree serves two purposes: (1) under a specified set of assumptions, it establishes an upper bound on performance, and (2) it discloses what characteristics of the tree structure are correlated with optimal performance.
V-A. Optimal Definition
We previously showed that CPU costs are the bottleneck that affect device performance. Intuitively, an optimal hash tree must therefore be a tree that reduces the number of hashes that must be computed per update or verification, reducing hashing costs and therein improving performance.